Please help me solve this equation: 1-3c=9c+7
Found 3 solutions by sparks211a, unlockmath, Alan3354:
Answer by sparks211a(3)

1-3c=9c+7
-3c=9c+6
0=12c+6
-6=12c
Divide both sides by 12
c=-6/12
c = -1/2

Answer by unlockmath(1688)

Hello,
The key to solving an equation is getting the variable (in this case, "c") all by itself on one side of the equal sign.
1-3c=9c+7 Add 3c to both sides to get:
1=12c+7 Now subtract 7 from both sides to get:
-6=12c Now the last step is to divide both sides by 12 to result in:
c=-1/2
Plug this answer into the original to see if it works.
RJ
There's a book I wrote that will help you out.
www.math-unlock.com

Answer by Alan3354(69443)

1-3c=9c+7
1 = 12c + 7
-12c = 6
c = 6/-12
c = -1/2